Solve the problem by stating the equation you would use to solve and then solving it:
after spending $3850 for tables and $1750 for chairs a convention center manager finds that 30% of his original budget remains. find the amount that
remains. round to the nearst dollar.

Answer by ExplanationCentral(10)

3850 + 1750 = 5600.
So the problem tells us that this 5600 would be 70% of his original budget.
Let the original budget be x. 5600 = 0.7(x). So x = 5600/0.7 = 8000.
If 30% of his budget remains, this is 0.3*8000 = 2400.